OpenStudy (anonymous):

squarerrot(1-cosx/1+cosx)=cosecx-cotx

OpenStudy (anonymous):

\[LHS=\sqrt{\frac{1-cosx}{1+cosx}}=\sqrt{\frac{1-cosx}{1+cosx}*\frac{1-cosx}{1-cosx}}\] \[=\sqrt{\frac{(1-cosx)^2}{1-\cos^2x}}\] \[=\sqrt{\frac{(1-cosx)^2}{\sin^2x}}\] \[=\frac{1-cosx}{sinx}\] \[=\frac{1}{sinx}-\frac{cosx}{sinx}\] \[=cosecx-cotx=RHS\]

OpenStudy (anonymous):

prove??

OpenStudy (anonymous):

yes thanks